OpenStudy (anonymous):

What is the quickest way to solve for X_3(s)/F(s) in a 3x3 matrix can someone help me?

OpenStudy (anonymous):

\[\left(\begin{matrix} X_1(s) \\X_2(s) \\ X_3(s) \end{matrix}\right) = \left[\begin{matrix}a & b & c\\ d & e & f\ \\ g & h & i \end{matrix}\right]\left(\begin{matrix}0 \\ F(s) \\ 0 \end{matrix}\right)\]

OpenStudy (anonymous):

could I do (g*0+hF(s) + i*0 )/determinant?

OpenStudy (loser66):

@e.mccormick your field. right?

OpenStudy (loser66):

@allsmiles I really don't understand your question, can you please post the original problem?

OpenStudy (anonymous):

I wrote the matrix out wrong sorry. But my question is to just solve for X_3/F(s) give this matrix equation

OpenStudy (anonymous):

I will rewrite the matrix

OpenStudy (anonymous):

\[\left(\begin{matrix}X_1 \\ X_2 \\ X_3 \end{matrix}\right) = \left[\begin{matrix}a & b & c \\ d & e & f \\ g & h & i\end{matrix}\right]^{-1}\left(\begin{matrix}0 \\ F(s) \\ 0\end{matrix}\right) \]

OpenStudy (anonymous):

F(s) being a function. I want to solve for X_3 / F(s)

OpenStudy (e.mccormick):

You have an inverse the second time, not the first...

OpenStudy (anonymous):

Yeah I wrote my matrix out wrong I apologize

OpenStudy (e.mccormick):

kk. So this was original;ly Ax=b, and you did A^{-1}Ax=A^{-1}b to get it to that point?

OpenStudy (anonymous):

yes

OpenStudy (e.mccormick):

It would depend on it being an invertable matrix for that to work... but if you do that, then it is just a matrix multiplication at that point.

OpenStudy (anonymous):

ah sorry I don't quite understand my book got this

OpenStudy (anonymous):

\[X_3 = \left[\begin{matrix}a & b & 0 \\ d & e & F(s) \\ g & h & 0\end{matrix}\right]/\left[\begin{matrix}a & b & c \\ d & e & f \\ g & h & i \end{matrix}\right]\]

OpenStudy (anonymous):

how did they get to this point? I can solve it that way

OpenStudy (e.mccormick):

Is the second part of that [] or | |. THey have different meanings.

OpenStudy (anonymous):

oh they are both | | sorry, does that stand for finding the determinant

OpenStudy (e.mccormick):

Yes.

OpenStudy (anonymous):

yeah both matrix are | |. so it's | | / | |

OpenStudy (anonymous):

but I have no idea how they got that expression

OpenStudy (e.mccormick):

Through doing the assorted operations. Seen this: \(A^{-1}=\frac{1}{|A|}adj(A)\)

OpenStudy (anonymous):

but then why is the expression equal to just X_3?

OpenStudy (e.mccormick):

Beacuse once you find the adjoint, it is multiplied by that vector.

OpenStudy (anonymous):

I am quite new to matrices, does adjoint just mean inverse?

OpenStudy (e.mccormick):

No. the inverse of the determinant times the adjoint is the inverse. The adjoint is the transpose of the cofactors and the cofactors is the minors with signs applied.

OpenStudy (anonymous):

Oh I see okay so the top matrix of the | | / | | I posted above is finding the adjoint of A and the bottom is the determinant of A?

OpenStudy (anonymous):

I'm still kind of lost on how they replaced the right side with 0 F(s) 0

OpenStudy (anonymous):

sorry for so many questions..

OpenStudy (e.mccormick):

Well, they said it was that, right? Did it actually start with this: \[\left[\begin{matrix}a & b & c \\ d & e & f \\ g & h & i\end{matrix}\right]\left(\begin{matrix}X_1 \\ X_2 \\ X_3 \end{matrix}\right)=\left(\begin{matrix}0 \\ F(s) \\ 0\end{matrix}\right)\]

OpenStudy (anonymous):

Yeah that's how it started

OpenStudy (e.mccormick):

Now, do you know what that right hand side would look like if you multiplied it?

OpenStudy (e.mccormick):

Sorry... left hand side....

OpenStudy (anonymous):

woudln't you get 3 equations..

OpenStudy (anonymous):

lol oops I meant aX_1 + bX_2+cX_3 = 0

OpenStudy (anonymous):

so if I want to solve

OpenStudy (anonymous):

for the X_2 I would replace the second column by 0 F(s) 0 instead

OpenStudy (e.mccormick):

Yah, looks like the email worked... lets see if I can post again finally.

OpenStudy (anonymous):

okay should I still go about finding the inverse?

OpenStudy (e.mccormick):

According to Craimer's Rule, if you have any matrix A, and vectors x and b, if Ax=b, you can form new matrices by replacing columns with the b. For this, lets take a 3x3 because that is what you have. \(A_1\) is A but with column 1 replaced by b. \(A_2\) has column 2 replaced, and so on. The the corresponding x values are the determinant of the new matrix divided by the determinant of the original, or precisely that: \[\left[\begin{matrix}a & b & c \\ d & e & f \\ g & h & i\end{matrix}\right]\left(\begin{matrix}X_1 \\ X_2 \\ X_3 \end{matrix}\right)=\left(\begin{matrix}0 \\ F(s) \\ 0\end{matrix}\right)\implies \\ X_3 = \frac{\left|\begin{matrix}a & b & 0 \\ d & e & F(s) \\ g & h & 0\end{matrix}\right|}{\left|\begin{matrix}a & b & c \\ d & e & f \\ g & h & i \end{matrix}\right|}\]With the between step being just making \(A_3\) to put on the top.

OpenStudy (e.mccormick):

No point to finding the inverse for that.... I mean, it could be done, but it is a ton of math for no reason.

OpenStudy (e.mccormick):

Did you need it solved past that point? If so, do cofactor expansion on the third column.

OpenStudy (anonymous):

Oh snap thanks so much I get it!!!

OpenStudy (anonymous):

but what rule did you know to solve for X_3?

OpenStudy (anonymous):

like dividing it by the determinant?

OpenStudy (e.mccormick):

Oh, I could use any of them. That is Craimer's Rule.

OpenStudy (anonymous):

Oh that is so cool does this work with any size of matrices?

OpenStudy (anonymous):

Thanks so much for your help and writing all that out I really appreciate it!!

OpenStudy (e.mccormick):

By the same rule: \[X_2 = \frac{\left|\begin{matrix}a & 0 & c \\ d & F(s) & f \\ g & 0 & i\end{matrix}\right|}{\left|\begin{matrix}a & b & c \\ d & e & f \\ g & h & i \end{matrix}\right|}\\ X_1 = \frac{\left|\begin{matrix}0 & b & c \\ F(s) & e & f \\ 0 & h & i\end{matrix}\right|}{\left|\begin{matrix}a & b & c \\ d & e & f \\ g & h & i \end{matrix}\right|}\] Craimer's rule is valid for any size matrix. If you actually take those determinants, you can solve it a little bit more. Don't know if there would be any reason to with this.

OpenStudy (e.mccormick):

See how which column I replaces is in line with what column it would be if I multiplied out the original equation? All the \(X_1\) values are replaced with the solution when solving for \(X_1\).

OpenStudy (anonymous):

Yes I see how it works now, wish I could give you more stars. You have no idea how much this will help me, thanks again!

OpenStudy (e.mccormick):

And do you know what cofactor expansion is? Have they gone over that?